# Weak interaction axial form factors of the octet baryons in nuclear medium

### v1

@@ Abstract
We study the axial-vector and the induced pseudoscalar form factors associated with the weak transitions between the octet baryon members in nuclear medium, using a covariant constituent quark model. We extend previous calculations of the axial transition form factors from the vacuum (free space) to the nuclear medium (symmetric nuclear matter). The extension of the model to the nuclear medium takes into account the modifications of the properties of hadrons in the medium (masses and coupling constants), as determined by the quark-meson coupling model. The axial-vector ($G_{A}$) and the induced pseudoscalar ($G_{P}$) form factors are evaluated for different values of the nuclear density $\rho$ in terms of the square transfer momentum $q^{2}=-Q^{2}$. We conclude that in general the $G_{A}$ and $G_{P}$ form factors are reduced in the nuclear medium. The reduction is stronger for light baryons and high densities. The medium modifications are milder for the heavier octet baryons, particularly at large $Q^{2}$. The calculations presented here can be used to estimate the cross sections of neutrino and antineutrino scattering with nucleus, and neutrino and antineutrino scattering with hyperons bound to a nucleus, as well as those in the cores of compact stars.
preprint: LFTC-24-06/89
@@ I Introduction
In the recent years a considerable effort has been made in the experimental and theoretical studies of the internal structure of hadrons, including the electromagnetic and weak structures of the nucleon and other baryons in the vacuum [1, 2, 3, 7, 4, 5, 6, 10, 1, 8, 9, 11], and in nuclear matter, and dense nuclear medium [13, 14, 15, 16, 17, 18, 19, 20, 21, 12]. The normal nuclear matter may be characterized by the saturation density $\rho_{0}\simeq 0.15$ fm-3, and the energy per nucleon $\simeq-15.7$ MeV. The structure of the hadrons is modified when immersed in the nuclear medium. Studies in vacuum are important to set a reference to study in nuclear matter. Studies in the nuclear medium are important for the understanding of environments with dense nuclear matter from the high energy nucleus-nucleus collisions to the cores of compact stars (neutron stars and magnetars) [28, 23, 25, 26, 24, 27, 22].
The present work highlights on the study of the axial structure of baryons in nuclear matter. The main motivation is to develop a formalism that can be used to make predictions of physical properties of baryons at relatively high densities as well as moderate, and large $Q^{2}$, based on the degrees of freedom manifest in the free space: valence quarks and meson cloud excitations of the baryon cores. Within the formalism one can estimate the modifications of the axial transition form factors due to the nuclear medium, by comparison with the corresponding functions in the free space [28, 29]. Deviations between the in-medium and in-vacuum form factors are a manifestation of the role of the medium on the internal structure of the baryons. The possible effects were observed already in electromagnetic reactions [21, 30, 31, 32], although the interpretations of the observed modifications in connection with the European Muon Collaboration (EMC) effect [33] are still in debate [34, 35]. The calculations presented here can be used in the study of reactions of neutrinos/antineutrinos with nucleus and reactions of neutrinos/antineutrinos with hyperons in dense nuclear matter. The methods developed here for the octet baryons can be extended in the future to the other baryon systems, like decuplet baryons and transitions between octet baryons and decuplet baryons, as well as for densities larger than the normal nuclear matter.
@@
Although there are not much recent developments from the experimental side, there has been a significant progress in lattice QCD simulations for weak axial-vector couplings and transition form factors of the nucleon [36, 37, 38, 39, 40, 41, 42, 43, 44], as well as for the axial-vector couplings of the octet baryons [45, 46, 47]. Lattice QCD simulations for pion mass values close to the physical limit reproduce accurately the experimental value of the nucleon axial-vector coupling, $g_{A}=G_{A}(0)$, but there are still discrepancies on the axial-vector form factors $G_{A}(Q^{2})$ for finite $Q^{2}$ [36, 37].
The combined analysis of lattice simulations and experimental data for large $Q^{2}$, with model calculations based on different frameworks, suggest that the octet baryon weak axial-vector form factors can be interpreted as the result of a dominant contribution associated with the valence quark and a component associated with the meson cloud dressing of the baryon cores [8, 48].
@@
The model based on the SU(6) spin-flavor symmetry for the quark states has been applied in particular, in the study of the octet baryons, the decuplet baryons and transitions between the octet and decuplet baryons [9, 28, 29, 58, 59, 60, 61, 63, 62, 64]. The starting point is the representation of the quark weak axial structure in terms of quark axial form factors parametrized in a vector meson dominance form. The free parameters of the models are fixed by the study of the weak axial form factors ($G_{A}$ and $G_{P}$) in the lattice QCD regime for the nucleon. Once calibrated the model (quark form factors and radial wave functions), the calculations are extrapolated for the octet baryons in the physical regime. The contributions of the valence quarks are combined with a phenomenological SU(3) parametrization of the meson cloud contribution. The comparison between model estimates for the nucleon and the experimental data are used to estimate the magnitude of the contributions from the meson cloud effects.
The model for the vacuum is in the present work extended to the study of the octet baryon weak axial form factors in symmetric nuclear matter. The methodology used is based on the extension of a model developed previously in vacuum [8] to the nuclear medium. The extension uses the quark-meson coupling (QMC) model [23], to estimate the masses of the baryons and the mesons, as well as the baryon-meson couplings in nuclear medium. The valence quark component of the model in medium is obtained considering the natural generalization of the covariant spectator quark model based on the vector meson dominance parametrization of the quark current and radial wave functions expressed in terms of the mass of the baryons [28]. The extension of meson cloud component is performed taking into account the medium modifications of the baryon-meson couplings. The weak axial form factors are then calculated in nuclear medium for different values of the nuclear density $\rho$.
Studies of the axial structure of the octet baryons based on different frameworks can be found in Refs. [65, 68, 66, 67] for the vacuum, and in Refs. [69, 70, 71] for nuclear matter.
@@ II.1 Axial-vector current
where $u_{B^{\prime}}$ and $u_{B}$ are the corresponding Dirac spinors, and $M_{BB^{\prime}}$ is the average mass of the final (mass $M_{B^{\prime}}$) and initial (mass $M_{B}$) baryons $M_{BB^{\prime}}=\frac{1}{2}(M_{B^{\prime}}+M_{B})$. The factor $1/2$ in Eq. (1) is included to be consistent with the nucleon case [8]. The form factors $G_{A}$ and $G_{P}$ depend then on the indices $B$ and $B^{\prime}$. For simplicity we omit those indices when possible.
The expression Eq. (1) is obtained after a proper projection in the flavor space. Those flavor operators can be represented in terms of the Gell-Mann matrices $\lambda_{i}$ ($i=1,..,8$[5] as discussed in Appendix. For the present study the cases of interest are the operators: $I_{0}=\lambda_{3}$ (neutral transitions), $I_{\pm}$ (increases/decreases the isospin projection) and $V_{\pm}$ (decreases/increases the number of strange quarks).
@@ II.2 Experimental status
We start with the case of the nucleon. The function $G_{A}$ can be measured by neutrino scattering and pion electroproduction off nucleons. Both experiments suggest a dipole type function dependence $G_{A}(Q^{2})=G_{A}(0)/(1+Q^{2}/M_{A}^{2})^{2}$, where the values of $M_{A}$ varies between 1.03 and 1.07 GeV depending on the method [4, 73].
@@
There are also nucleon data for $G_{P}$ in the range $Q^{2}=$0–0.2 GeV2 determined by muon capture [4] and the from pion electroproduction data using the low energy theorem [76].
Concerning the other members of the octet baryons, there are data for $G_{A}(0)$ determined by the semileptonic decays indicated in Table 1. The possible transitions between the octet baryon members are discussed in the next section.
@@ II.3 Theory
where $\left|qqq\right>$ is the three-quark state and $c_{B}\left|{\rm MC}\right>$ represents the baryon-meson state associated with the meson cloud dressing. The coefficient $c_{B}$ is determined by the normalization $Z_{B}(1+c_{B}^{2})=1$, assuming that $\left|{\rm MC}\right>$ is normalized to the unity.
@@
In Eq. (3), we include only the first order, or ”one meson in the air” correction for the meson cloud, associated with the baryon-meson states. In principle, we should also include corrections associated with baryon-meson-meson states. These corrections are in general small due to the suppression associated with the multiple meson excitations. One can exemplify this effect using the nucleon case as example. The nucleon meson cloud is dominated by the $|N\pi\rangle$ with $N\equiv qqq$ component. Based on the previous work [8], we can estimate the probability of the $|N\pi\rangle$ state as $1-Z_{N}\approx 0.3$. The correction associated with the two pion plus $N=qqq$ state ($|N\pi\pi\rangle$) contribution is then attenuated by the factor $(1-Z_{N})^{2}\approx 0.09$.
@@ III Model for the vacuum
The covariant spectator quark model is a constituent quark model derived from the covariant spectator theory [49, 7, 77]. Within the framework, the baryons are described as three-constituent quark systems, where a quark is free to interact with electroweak fields in the impulse approximation. The electromagnetic interaction with the quarks is described in terms of quark electromagnetic and axial form factors, that simulate the structure associated with the gluon and quark-antiquark dressing of the quarks. The constituent quark form factors are parametrized using a vector meson dominance (VMD) picture calibrated in the study of the electromagnetic structure of the nucleon and baryon decuplet [49, 50].
In the electroweak interaction with the baryon systems the probes act on the individual quarks in the relativistic impulse approximation. In the covariant spectator theory formalism [77] one can then integrate over the quark-pair degrees of freedom since they are spectators in the interaction, and reduce the system to a diquark-quark state, where the diquark can be represented as an on-mass-shell particle with effective mass $m_{D}$ [49, 51]. The baryon wave functions are derived from spin-flavor-radial symmetries associated with the diquark-quark configurations. The radial wave functions are determined phenomenologically by experimental data, or lattice QCD results for some ground state systems [49, 50, 54, 9, 28, 60, 55].
@@
$G_{P}^{\rm Pole}=\frac{4M_{BB^{\prime}}^{2}}{\mu^{2}+Q^{2}}\tilde{G}_{A}^{\rm B},$ (6)
where $\mu=m_{\pi}$ (pion mass) for $\Delta I=1$ transitions and $\mu=m_{K}$ (kaon mass) for $\Delta S=1$ transitions, as discussed in Ref. [8].
@@ III.1 Bare contribution
where $n_{P}$ is the $P$-state admixture coefficient and $n_{S}=\sqrt{1-n_{S}^{2}}$.
@@
where $\phi_{S}^{0,1}$ are the spin wave functions labeled by the diquark spin ($S=0,1$), $\left|M_{S,A}\right>$ are the flavor states, symmetric ($S$) or anti-symmetric ($A$) in the exchange of the flavors between the quarks 1 and 2, $\psi_{S}$ and $\psi_{P}$ are the radial wave functions and $\tilde{k}=k-\frac{P\cdot k}{M^{2}}P$. The flavor states $\left|M_{S,A}\right>$ for the octet baryons are presented in Table 2.
@@
$j_{Aq}^{\mu}=\left(g_{A}^{q}(Q^{2})\gamma^{\mu}+g_{P}^{q}(Q^{2})\frac{q^{\mu}}{2M}\right)\gamma_{5}\frac{\lambda_{a}}{2},$ (10)
where $\lambda_{a}$ ($a=1,...,8$) are the Gell-Mann matrices and $g_{A}^{q}$ and $g_{P}^{q}$ are the quark axial-vector and induced pseudoscalar form factors, respectively. In Eq. (10) the flavor operators $\lambda_{a}$ act on the quark flavor states ($\left|M_{A,S}\right>$) and the Lorentz operators on the baryon spin states. The explicit form of the matrices $\lambda_{a}$ are presented in Appendix.
@@
where $l=2$ ($N$), $l=3$ ($\Lambda$ and $\Sigma$) or $l=4$ ($\Xi$). Numerically we use the parametrization from Ref. [28]. Notice that $\beta_{2}>\beta_{3}>\beta_{4}$ reproduce the natural size of the baryons, where systems with more strange quarks are more compact. The radial wave functions for the $P$-states are related to the $S$-state radial wave functions [8].
In the covariant spectator quark model the weak transition form factors are determined in the relativistic impulse approximation by the current
@@
where the index $a$ in Eq. (10) is determined by the flavors associated with the transition $B\to B^{\prime}$.
@@
$\Delta I=1\hskip 10.00002pt$ ($I_{+}$)       $n\to p$ $1$ $-\frac{1}{3}$        $\frac{10}{9}$
@@
$\Delta S=1\hskip 10.00002pt$ ($V_{+}$)       $\Lambda\to p$ $-\frac{2}{\sqrt{6}}$ 00        $-\sqrt{\frac{2}{3}}$
@@
The results obtained for the form factors for the wave functions (7) and current (10), are
@@
The flavor transition operators are discussed in Appendix. The coefficients $f_{X}^{A,S}$ and the factor ${\cal F}$ are presented in Table 3.
In the previous equations $\tilde{G}_{A}$, $\tilde{G}_{P}$, $g_{A}^{q}$, $g_{P}^{q}$ and $B_{i}$ are functions of $Q^{2}$. The explicit expressions for $B_{i}$ ($i=0,...,5$) can be found in Ref. [8]. The normalization of the radial wave functions lead to $B_{0}(0)=1$. The effect of the mass ($M$ or $M_{BB^{\prime}}$) on the transition form factors appears on the functions $B_{i}$.
The leading order contributions for $\tilde{G}_{A}^{\rm B}$ and $\tilde{G}_{P}^{\rm B}$ comes from the $S$-state contribution. In the $S$-state limit ($n_{P}=n_{SP}=0$) the form factors are proportional to $\frac{3}{2}{\cal F}$ and we recover nonrelativistic results. For the nucleon $\tilde{G}_{A}^{\rm B}\simeq\frac{5}{3}g_{A}^{q}B_{0}$. Different SU(3) coefficients are obtained for the other transitions.
In Table 3, the transitions $\Delta I=1$ correspond to $d\to u$ transition, except for the $\Sigma^{+}\to\Lambda$ ($u\to d$ transition). From the table, we can conclude that the $\Sigma^{+}\to\Lambda$ and $\Sigma^{-}\to\Lambda$ transition form factors differ by a sign, and that the $\Sigma^{0}\to\Sigma^{+}$ and $\Sigma^{-}\to\Sigma^{+}$ transition form factors are identical. One can then conclude that in the case $\Delta I=1$ there are only four cases of interest.

### v2

@@ Abstract
We study the axial-vector and the induced pseudoscalar form factors associated with the weak transitions between the octet baryon members in nuclear medium, using a covariant constituent quark model. We extend previous calculations of the axial transition form factors from the vacuum (free space) to the nuclear medium (symmetric nuclear matter). The extension of the model to the nuclear medium takes into account the modifications of the properties of hadrons in the medium (masses and coupling constants), as determined by the quark-meson coupling model. The axial-vector ($G_{A}$) and the induced pseudoscalar ($G_{P}$) form factors are evaluated for different values of the nuclear density $\rho$ in terms of the square transfer momentum $q^{2}=-Q^{2}$. We conclude that, in general, the $G_{A}$ and $G_{P}$ form factors are reduced in the nuclear medium. The reduction is stronger for light baryons and high densities. The medium modifications are milder for the heavier octet baryons, particularly at large $Q^{2}$. The calculations presented here can be used to estimate the cross sections of neutrino and antineutrino scattering with nucleus, and neutrino and antineutrino scattering with hyperons bound to a nucleus, as well as those in the cores of compact stars.
@@ I Introduction
In recent years a considerable effort has been made in the experimental and theoretical studies of the internal structure of hadrons, including the electromagnetic and weak structures of the nucleon and other baryons in the vacuum [1, 2, 3, 8, 4, 5, 6, 10, 1, 7, 9, 11], and in nuclear matter, and dense nuclear medium [13, 14, 15, 16, 17, 18, 19, 20, 21, 12]. The normal nuclear matter may be characterized by the saturation density $\rho_{0}\simeq 0.15$ fm-3, and the energy per nucleon $\simeq-15.7$ MeV. The structure of the hadrons is modified when immersed in the nuclear medium. Studies in vacuum are important to set a reference to study in nuclear matter. Studies in the nuclear medium are important for the understanding of environments with dense nuclear matter from the high-energy nucleus-nucleus collisions to the cores of compact stars (neutron stars and magnetars) [28, 23, 25, 26, 24, 27, 22]. At the moment, however, there are not many theoretical studies about the internal structure modifications of baryons in nuclear matter.
The present work highlights the study of the axial structure of baryons in nuclear matter. The main motivation is to develop a formalism that can be used to make predictions of physical properties of baryons at relatively high densities as well as moderate, and large $Q^{2}$, based on the degrees of freedom manifest in the free space: valence quarks and meson cloud excitations of the baryon cores. Within the formalism one can estimate the modifications of the axial transition form factors due to the nuclear medium, by comparison with the corresponding functions in the free space [28, 29]. Deviations between the in-medium and in-vacuum form factors are a manifestation of the role of the medium on the internal structure of the baryons. The possible effects were observed already in electromagnetic reactions [21, 30, 31, 32], although the interpretations of the observed modifications in connection with the European Muon Collaboration (EMC) effect [33] are still in debate [34, 35]. The calculations presented here can be used in the study of reactions of neutrinos/antineutrinos with nucleus and reactions of neutrinos/antineutrinos with hyperons in dense nuclear matter. The methods developed here for the octet baryons can be extended in the future to the other baryon systems, like decuplet baryons and transitions between octet baryons and decuplet baryons, as well as for densities larger than the normal nuclear matter.
@@
Although there are not many recent developments from the experimental side, there has been a significant progress in lattice QCD simulations for weak axial-vector couplings and transition form factors of the nucleon [36, 37, 38, 39, 40, 41, 42, 43, 44], as well as for the axial-vector couplings of the octet baryons [45, 46, 47]. Lattice QCD simulations for pion mass values close to the physical limit reproduce accurately the experimental value of the nucleon axial-vector coupling, $g_{A}=G_{A}(0)$, but there are still discrepancies on the axial-vector form factors $G_{A}(Q^{2})$ for finite $Q^{2}$ [36, 37].
The combined analysis of lattice simulations and experimental data for large $Q^{2}$, with model calculations based on different frameworks, suggests that the octet baryon weak axial-vector form factors can be interpreted as the result of a dominant contribution associated with the valence quark and a component associated with the meson cloud dressing of the baryon cores [7, 48].
@@
The model based on the SU(6) spin-flavor symmetry for the quark states has been applied in particular, in the study of the octet baryons, the decuplet baryons and transitions between the octet and decuplet baryons [9, 28, 29, 58, 59, 60, 61, 63, 62, 64]. The starting point is the representation of the quark weak axial structure in terms of quark axial form factors parametrized in a vector meson dominance form. The free parameters of the models are fixed by the study of the weak axial form factors ($G_{A}$ and $G_{P}$) in the lattice QCD regime for the nucleon. Once the model is calibrated (quark form factors and radial wave functions), the calculations are extrapolated for the octet baryons in the physical regime. The contributions of the valence quarks are combined with a phenomenological SU(3) parametrization of the meson cloud contribution. The comparison between model estimates for the nucleon and the experimental data is used to estimate the magnitude of the contributions from the meson cloud effects.
The model for the vacuum is in the present work extended to the study of the octet baryon weak axial form factors in symmetric nuclear matter. The methodology used is based on the extension of a model developed previously in vacuum [7] to the nuclear medium. The extension uses the quark-meson coupling (QMC) model [23], to estimate the masses of the baryons and the mesons, as well as the baryon-meson couplings in nuclear medium. The QMC model extends the properties of the bag model [23, 65, 66] to the nuclear matter. The cloudy bag model has already been combined with the covariant spectator quark model in the study of the electromagnetic transitions between baryon states in vacuum [62, 63, 61]. In these works the correspondence between the quark effective electromagnetic structure in the two models is used to calculate the electromagnetic interaction with baryons in intermediate states. The valence quark component of the model in medium is obtained considering the natural generalization of the covariant spectator quark model based on the vector meson dominance parametrization of the quark current and radial wave functions expressed in terms of the mass of the baryons [28]. The extension of the meson cloud component is performed using an effective SU(3) baryon-meson parametrization [7] and taking into account the medium modifications of the baryon-meson couplings. The weak axial form factors are then calculated in nuclear medium for different values of the nuclear density $\rho$.
The first studies of the electromagnetic and axial structure of baryons were made using Skyrme and soliton models [67, 68, 69, 70, 71, 72, 73, 80, 79, 81, 74, 75, 76, 78, 77] and the QMC model [23, 27, 24, 25, 26]. Other studies of the axial structure of the nucleon and octet baryons based on different frameworks in vacuum can be found in Refs. [82, 84, 83, 85, 86, 87, 88, 90, 89, 91, 92, 93, 94, 95, 96, 97], while, studies of the electromagnetic and axial structure of the nucleon and octet baryons in medium can be found in Refs. [101, 98, 99, 100, 102, 104, 105, 106, 107, 108, 109, 103, 111, 110, 112].
@@ II.1 Axial-vector current
where $u_{B^{\prime}}$ and $u_{B}$ are the corresponding Dirac spinors, and $M_{BB^{\prime}}$ is the average mass of the final (mass $M_{B^{\prime}}$) and initial (mass $M_{B}$) baryons $M_{BB^{\prime}}=\frac{1}{2}(M_{B^{\prime}}+M_{B})$. The factor $1/2$ in Eq. (1) is included to be consistent with the nucleon case [7]. The form factors $G_{A}$ and $G_{P}$ depend then on the indices $B$ and $B^{\prime}$. For simplicity, we omit those indices when possible.
The expression Eq. (1) is obtained after a proper projection in the flavor space. Those flavor operators can be represented in terms of the Gell-Mann matrices $\lambda_{i}$ ($i=1,..,8$[5] as discussed in Appendix A. For the present study the cases of interest are the operators $I_{0}=\lambda_{3}$ (neutral transitions), $I_{\pm}$ (increases/decreases the isospin projection) and $V_{\pm}$ (decreases/increases the number of strange quarks).
@@ II.2 Experimental status
We start with the case of the nucleon. The function $G_{A}$ can be measured by neutrino scattering and pion electroproduction off nucleons. Both experiments suggest a dipole-type function dependence $G_{A}(Q^{2})=G_{A}(0)/(1+Q^{2}/M_{A}^{2})^{2}$, where the value of $M_{A}$ varies between 1.03 and 1.07 GeV depending on the method [4, 113].
@@
There are also nucleon data for $G_{P}$ in the range $Q^{2}=$0–0.2 GeV2 determined by muon capture [4] and from pion electroproduction data using the low-energy theorem [116].
Concerning the other members of the baryon octet, there are data for $G_{A}(0)$ determined by the semileptonic decays indicated in Table 1. The possible transitions between the octet baryon members are discussed in the next section.
@@ II.3 Theory
where $\left|qqq\right>$ is the three-quark state and $c_{B}\left|{\rm MC}\right>$ represents the baryon-meson state associated with the meson cloud dressing. The coefficient $c_{B}$ is determined by the normalization $Z_{B}(1+c_{B}^{2})=1$, assuming that $\left|{\rm MC}\right>$ is normalized to unity.
@@
In Eq. (3), we include only the first order, or ”one meson in the air” correction for the meson cloud, associated with the baryon-meson states. In principle, we should also include corrections associated with baryon-meson-meson states. These corrections are, in general, small due to the suppression associated with the multiple meson excitations. One can exemplify this effect using the nucleon case as an example. The nucleon meson cloud is dominated by the $|N\pi\rangle$ with $N\equiv qqq$ component. Based on previous work [7], we can estimate the probability of the $|N\pi\rangle$ state as $1-Z_{N}\approx 0.3$. The correction associated with the two pion plus $N=qqq$ state ($|N\pi\pi\rangle$) contribution is then attenuated by the factor $(1-Z_{N})^{2}\approx 0.09$.
The decomposition (3) is model dependent since the contributions from the meson cloud cannot be disentangled in experiments. The model dependence is a consequence of the relation between the meson cloud contributions and the scale associated to the bare core [117, 118, 119]. In Sec. III.2, we discuss how the model dependence can be reduced by the comparison with lattice QCD simulations with large pion masses [8, 52, 61].
@@ III Model for the vacuum
The covariant spectator quark model is a constituent quark model derived from the covariant spectator theory [49, 8, 120]. Within the framework, the baryons are described as three-constituent quark systems, where a quark is free to interact with electroweak fields in the impulse approximation. The electromagnetic interaction with the quarks is described in terms of quark electromagnetic and axial form factors, that simulate the structure associated with the gluon and quark-antiquark dressing of the quarks. The constituent quark form factors are parametrized using a vector meson dominance picture calibrated in the study of the electromagnetic structure of the nucleon and baryon decuplet [49, 50].
[…]
@@
$G_{P}^{\rm Pole}=\frac{4M_{BB^{\prime}}^{2}}{\mu^{2}+Q^{2}}G_{A}^{\rm B},$ (6)
where $\mu=m_{\pi}$ (pion mass) for $|\Delta I|=1$ transitions and $\mu=m_{K}$ (kaon mass) for $|\Delta S|=1$ transitions, as discussed in Ref. [7].
@@ III.1 Bare contribution
where $P$ is the momentum of the baryon (mass $M_{B}$), $n_{P}$ is the $P$-state admixture coefficient and $n_{P}=\sqrt{1-n_{S}^{2}}$.
@@
where $\phi_{S}^{0,1}$ are the spin wave functions labeled by the diquark spin ($S=0,1$), $\left|M_{S,A}\right>$ are the flavor states, symmetric ($S$) or antisymmetric ($A$) in the exchange of the flavors between the quarks 1 and 2, $\psi_{S}$ and $\psi_{P}$ are the radial wave functions and $\tilde{k}=k-\frac{P\cdot k}{M_{B}^{2}}P$. The flavor states $\left|M_{S,A}\right>$ for the octet baryons are presented in Table 2.
@@
$j_{Aq}^{\mu}=\left(g_{A}^{q}(Q^{2})\gamma^{\mu}+g_{P}^{q}(Q^{2})\frac{q^{\mu}}{2M_{N}}\right)\gamma_{5}\frac{\lambda_{a}}{2},$ (10)
where $\lambda_{a}$ ($a=1,...,8$) are the Gell-Mann matrices and $g_{A}^{q}$ and $g_{P}^{q}$ are the quark axial-vector and induced pseudoscalar form factors, respectively. In Eq. (10) the flavor operators $\lambda_{a}$ act on the quark flavor states ($\left|M_{A,S}\right>$) and the Lorentz operators on the baryon spin states. The explicit form of the matrices $\lambda_{a}$ are presented in Appendix A. Notice in Eq. (10) the inclusion of the nucleon mass in the definition of the quark current. This property, used already in the definition of the electromagnetic current [49, 50] is fundamental for the extension of the model to the lattice QCD regime.
@@
where $l=2$ ($N$), $l=3$ ($\Lambda$ and $\Sigma$) or $l=4$ ($\Xi$). Numerically we use the parametrization from Ref. [28]. Notice that $\beta_{2}>\beta_{3}>\beta_{4}$ reproduce the natural size of the baryons, where systems with more strange quarks are more compact.
In the covariant spectator quark model the weak transition form factors are determined in the relativistic impulse approximation by the current [49, 50, 51]
@@
where the index $a$ in Eq. (10) is determined by the flavors associated with the transition $B\to B^{\prime}$. The integration symbol represent the integration on the diquark on-shell momentum.
In the previous equation we consider the axial coupling with the single quark (label 3) when the spectator quarks have labels 1 and 2 [diquark (12)]. The remaining contributions associated with the diquarks (13) and (23) are identical due to the symmetry in the change of the quark labels. The final contribution can then be obtained considering the contribution of a single diquark state multiplied by the factor 3 [49, 51]. The final expression for the transition current (15) based on a single quark operator (i.e. no exchange or interaction currents) is valid when we consider a phenomenological quark-diquark radial wave function with effective parameters determined by physical or lattice QCD data. A more detailed derivation of the relativistic impulse approximation within the covariant spectator quark model framework can be found in Ref. [51].
@@
With the form (14) chosen for the radial wave functions, including the factor $1/m_{D}$, the transition form factors are independent of the diquark mass [49, 50]. As mentioned already, the SU(3) symmetry breaking due to the mass of the strange quark is taken into account in the range parameters of the wave functions determined by the global fit to the octet baryon electromagnetic form factors [28].
@@
$|\Delta I|=1\hskip 10.00002pt$ ($I_{+}$)       $n\to p$ $1$ $-\frac{1}{3}$        $\frac{10}{9}$
@@
$|\Delta S|=1\hskip 10.00002pt$ ($V_{+}$)       $\Lambda\to p$ $-\frac{2}{\sqrt{6}}$ 00        $-\sqrt{\frac{2}{3}}$
@@
The results obtained for the form factors for the wave functions (7) and current (10) are
@@
The flavor transition operators are discussed in Appendix A. The coefficients $f_{X}^{A,S}$ and the factor ${\cal F}$ are presented in Table 3.
In the previous equations $\tilde{G}_{A}$, $\tilde{G}_{P}$, $g_{A}^{q}$, $g_{P}^{q}$ and $B_{i}$ are functions of $Q^{2}$. The explicit expressions for $B_{i}$ ($i=0,...,5$) can be found in Ref. [7]. The normalization of the radial wave functions leads to $B_{0}(0)=1$. The effect of the mass ($M_{N}$ or $M_{BB^{\prime}}$) on the transition form factors appears on the functions $B_{i}$. The dependence of the transition form factors on the mass of the baryons was studied in detail in Ref. [7], demonstrating the relevance of taking into account implicitly the effects of the SU(3) flavor-symmetry breaking.
The leading-order contributions for $\tilde{G}_{A}^{\rm B}$ and $\tilde{G}_{P}^{\rm B}$ comes from the $S$-state contribution. In the $S$-state limit ($n_{P}=n_{SP}=0$) the form factors are proportional to $\frac{3}{2}{\cal F}$ and we recover nonrelativistic results. For the nucleon $\tilde{G}_{A}^{\rm B}\simeq\frac{5}{3}g_{A}^{q}B_{0}$. Different SU(3) coefficients are obtained for the other transitions.
In Table 3, the transitions $|\Delta I|=1$ correspond to $d\to u$ transition, except for the $\Sigma^{+}\to\Lambda$ ($u\to d$ transition). From the table, we can conclude that the $\Sigma^{+}\to\Lambda$ and $\Sigma^{-}\to\Lambda$ transition form factors differ by a sign, and that the $\Sigma^{0}\to\Sigma^{+}$ and $\Sigma^{-}\to\Sigma^{+}$ transition form factors are identical. One can then conclude that in the case $|\Delta I|=1$ there are only four cases of interest.